OpenStudy (greenbannanabob):

Marcy plays a game with a spinner and a 6-sided number cube. The spinner is divided into 4 equal sections numbered 1 to 4. Marcy spins the spinner and rolls the number cube. She records both numbers. What is the probability that at least one of the numbers is a 4?

OpenStudy (greenbannanabob):

I added 1/4+1/6 to 5/12, but the answer key says the answer is 9/24. I don't understand why.

OpenStudy (greenbannanabob):

to get 5/12*

OpenStudy (welshfella):

lets look at the possible outcomes 1. spinner = 4 , cube = 4 OR 2. spinner = 4 cube not 4 OR 3. spinner not 4 , cube = 4

OpenStudy (greenbannanabob):

What about none of them getting a 4?

OpenStudy (welshfella):

1. as these are independent events we multiply So P (4,4) = 1/4 * 1/6 = 1/24

OpenStudy (welshfella):

we are only concerned with one 4 or above. do you follow the above?

OpenStudy (greenbannanabob):

I get it now. So there is a 1/24 probability that both are 4. Then there is a 3/24 probability of the cube being 4. Then there is a 5/24 probability of the spinner being a 4. Then, when all three are added together, the answer is 9/24! Is this right?

OpenStudy (welshfella):

yes thats correct

OpenStudy (greenbannanabob):

Thank you!

OpenStudy (welshfella):

with these compound events where you gave one event AND another you multiply when its OR - you add.

OpenStudy (welshfella):

* have not gave